OpenStudy (anonymous):

State how many imaginary and real zeros the function has. f(x) = x4 - 15x2 - 16

OpenStudy (anonymous):

solve \(u^2-15u-16=0\) and then replace \(u\) by \(x^2\)

OpenStudy (anonymous):

this one factors easily, so you should be in good shape

OpenStudy (anonymous):

but how do i know how many are real and imiginary

OpenStudy (anonymous):

That's why you need to solve it to find out!

OpenStudy (anonymous):

there are 2 imaginary right? and 2 real?

OpenStudy (anonymous):

Descartes' rule of signs probably won't hep here, huh?

OpenStudy (anonymous):

@CliffSedge no

OpenStudy (anonymous):

did you factor? i get \[(u+1)(u-16)=0\] so \[(x^2+1)(x^2-16)=0\]

OpenStudy (anonymous):

thast what i got too

OpenStudy (anonymous):

that means \(x^2=-1\) two complex zeros, namely \(i, -i\)

OpenStudy (anonymous):

ok thank you!!!!